# A Purely Data-Driven Adaptive Impedance Matching Method Robust to Parasitic Effects

### v1

@@ VI Conclusion
In this work, we have proposed a purely data-driven adaptive impedance matching method that is robust to parasitic effects. First, we have designed RECBM-Net, a DNN that maps the operating state of a practical L-network to its corresponding S-parameters. Then, we have formulated the impedance matching task via the trained surrogate model as a mathematical optimization problem, and introduced two numerical optimization strategies with different online computational overhead. Finally, to avoid repeated RECBM-Net inference during matching, we have proposed IMS-Net that directly predicts the optimal solution in a single forward pass. Simulation results have shown that RECBM-Net achieves exceptionally high prediction accuracy of the S-parameters, with the absolute error for each component being on the order of $10^{-4}$. In the simulated mismatched scenarios, SAPSO has achieved higher matching accuracy than AD-Adam, with the trade-off of substantially increased computational overhead. IMS-Net has significantly reduced the online inference overhead while maintaining excellent performance, lowering the reflection coefficient below 0.2 for $95\%$ of samples.

### v2

@@ VI Conclusion
In this work, we have proposed a data-driven adaptive impedance matching method that is robust to parasitic effects. First, we have designed RECBM-Net, a DNN that maps the operating state of a practical L-network to its corresponding S-parameters. Then, we have formulated the impedance matching task via the trained surrogate model as a mathematical optimization problem, and introduced two numerical optimization strategies with different online computational overheads. Finally, to avoid repeated RECBM-Net inference during matching, we have proposed IMS-Net that directly predicts the optimal solution in a single forward pass.
Simulation results have shown that RECBM-Net achieves exceptionally high prediction accuracy of the S-parameters, with a MAE of $6.98\times 10^{-5}$ and a MRE of $0.123\%$ across both the real and imaginary components of all S-parameter entries. For 9000 simulated mismatched scenarios, SAPSO has reduced the magnitude of the reflection coefficient below 0.2 for 95.92% of the samples but has incured substantial online computational overhead, averaging 2097 RECBM-Net inferences per sample. By exploiting gradient information, AD-Adam has significantly reduced the computational overhead to 285 inferences per sample, with a slight reduction in compliance to 93.42%. IMS-Net has achieved the best trade-off between accuracy and computational overhead, requiring only a single inference to determine the matching solution while maintaining excellent performance, with the reflection coefficient magnitude reduced below 0.2 for $95\%$ of samples.
In summary, all three matching solution determination strategies are necessary. SAPSO and AD-Adam serve as the respective baselines for matching accuracy and online computational overhead in comparison with IMS-Net. While IMS-Net offers the optimal trade-off between matching accuracy and computational overhead, it requires additional model training and testing. Therefore, when TTM constraints are critical, SAPSO and AD-Adam are more advantageous than IMS-Net, as both can be deployed immediately after training RECBM-Net. Under such constraints, SAPSO is suitable for impedance matching tasks that require high accuracy but are subject to low real-time constraints. A representative example is the compensation for impedance variations introduced during the manufacturing process after antenna design. For applications with rapidly varying antenna impedances that impose high real-time demands on the impedance matching process, AD-Adam is preferable owing to its low computational overhead. Finally, if the product development cycle allows for additional offline model training and testing, IMS-Net is the best choice as it provides matching accuracy comparable to SAPSO while achieving the lowest online computational overhead.